# Generative Uncertainty in Diffusion Models

### v1

@@ 1 Introduction
Diffusion (and flow-matching) models [Sohl-Dickstein et al., 2015, Song et al., 2020a, b, Lipman et al., 2022] have recently pushed the boundaries of generative modeling due to their strong theoretical underpinnings and scalability. Across various domains, they have enabled the generation of increasingly realistic samples [Rombach et al., 2022, Esser et al., 2024, Li et al., 2024]. Despite the impressive progress, state-of-the-art models can still generate low quality images that contain artefacts and fail to align with the provided conditioning information. This poses a challenge for deploying diffusion models, as it can lead to a poor user experience by requiring multiple generations to manually find an artefact-free sample.
@@
    We show that our generative uncertainty strongly outperforms previous uncertainty-based approaches for filtering out poor samples [Kou et al., 2024, De Vita and Belagiannis, 2025]. Additionally, we achieve competitive performance with non-uncertainty-based methods, such as realism score [Kynkäänniemi et al., 2019] and rarity score [Han et al., 2023] (Section 4.1), while also highlighting the complementary benefits of uncertainty (Appendix A.5).
@@
    We propose effective strategies to reduce the sampling overhead of Bayesian uncertainty (Section 4.2) and demonstrate the applicability of our framework beyond diffusion models by applying it to a (latent) flow matching model (Section A.7).
@@ 2.1 Generative Modeling
Modern deep generative models like variational autoencoders (VAEs) [Kingma, 2013], generative adversarial networks (GANs) [Goodfellow et al., 2014], and diffusion models differ in their exact probabilistic frameworks and training schemes, yet share a common sampling recipe: start with random noise and transform it into a new data sample [Tomczak, 2022]. Specifically, let $\boldsymbol{x}\in\mathcal{X}$ denote a data sample and $\boldsymbol{z}\in\mathcal{Z}$ an initial noise. A new sample is generated by:
@@
where $p(\mathbf{z})$ is an initial noise (prior) distribution, typically a standard Gaussian $\mathcal{N}(0,I)$, and $g_{\theta}:\mathcal{Z}\rightarrow\mathcal{X}$ is a generator function with model parameters $\theta\in\mathbb{R}^{P}$.
@@
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ and $\{\beta_{s}\}_{s=1}^{T}$ is a noise schedule chosen such that $\boldsymbol{x}_{T}\sim\mathcal{N}(0,I)$ (approximately). In the backward process, a denoising network $f_{\theta}$ is learned via a simplified regression objective (among various possible parameterizations, see Song et al. [2020b] or Karras et al. [2022]):
$\!\!\!\mathcal{L}(\theta;\mathcal{D})=\mathbb{E}_{t,\boldsymbol{x}_{0},\boldsymbol{\epsilon}}\!\left[\big{|}\big{|}f_{\theta}(\sqrt{\bar{\alpha}_{t}}\boldsymbol{x}_{0}\!+\!\sqrt{1\!-\!\bar{\alpha_{t}}}\boldsymbol{\epsilon},t)-\boldsymbol{\epsilon}\big{|}\big{|}_{2}^{2}\right]\>$ (1)
where $\mathcal{D}=\{\boldsymbol{x}_{n}\}_{n=1}^{N}$ denotes a training dataset of images. After training, diffusion models generate new samples via a generator function, $g_{\hat{\theta}}$, which consists of sequentially applying the learned denoiser, $f_{\hat{\theta}}$, and following specific transition rules from samplers such as DDPM [Ho et al., 2020] or DDIM [Song et al., 2020a].
@@ 2.2 Bayesian Deep Learning
Bayesian neural networks (BNNs) go beyond point predictions and allow for principled uncertainty quantification [Buntine and Weigend, 1991, Neal, 1995, Kendall and Gal, 2017, Jospin et al., 2022]. Let $h_{\psi}:\mathcal{X}\rightarrow\mathcal{Y}$ denote a predictive model with parameters $\psi\in\mathbb{R}^{O}$ and $\mathcal{D}=\{(\boldsymbol{x}_{n},\boldsymbol{y}_{n})\}_{i=1}^{N}$ denote training data. Instead of finding a single fixed set of parameters, $\hat{\psi}=\arg\max\mathcal{L}(\psi;\mathcal{D})$, that maximizes an objective function $\mathcal{L}$, BNNs specify a prior $p(\psi)$ over model parameters and define a likelihood $p(\boldsymbol{y}|h_{\psi}(\boldsymbol{x}))$, which together yield a posterior distribution via Bayes rule: $p(\psi|\mathcal{D})\propto p(\psi)\prod_{n=1}^{N}p(\boldsymbol{y}_{n}|h_{\psi}(\boldsymbol{x}_{n}))$. Under this Bayesian view, a predictive model for a new test point $\boldsymbol{x}_{*}$ is then obtained via the posterior predictive distribution [Murphy, 2022]:
@@
For large models, finding the exact posterior distribution is computationally intractable, hence an approximate posterior $q(\psi|\mathcal{D})$ is used instead. Popular approaches for approximate inference include deep ensembles [Lakshminarayanan et al., 2017], variational inference [Blundell et al., 2015, Zhang et al., 2018], SWAG [Mandt et al., 2017, Maddox et al., 2019], and Laplace approximation [Daxberger et al., 2021a]. Moreover, to alleviate computational overhead, it is common to give a ‘Bayesian treatment’ only to a subset of parameters [Kristiadi et al., 2020, Daxberger et al., 2021b, Sharma et al., 2023]. Finally, the intractable expectation integral in the posterior predictive is approximated via Monte-Carlo (MC) sampling:
@@ 3 Generative Uncertainty via Bayesian Inference
Figure 1: Illustration of how we compute generative uncertainty for each random noise $\boldsymbol{z}$. For a generative model $g_{\theta}$, we sample $M$ sets of model parameters from our posterior distribution $q(\theta|\mathcal{D})$ and generate $M$ images. Then, we evaluate the semantic likelihood for each by computing feature embeddings with a pretrained encoder $c_{\phi}$ (e.g., CLIP) and take the uncertainty (e.g., entropy) over these embeddings. Random noises $\boldsymbol{z}$ with low uncertainty (left) tend to lead to consistent, high quality generations while random noises with high uncertainty (right) lead to poor, discordant generations.
@@ 3.1 Generative Uncertainty
Here, we use $\boldsymbol{z}$ (with a slight abuse of notation) to denote the entire randomness involved in the diffusion sampling process. [^1] In DDIM [Song et al., 2020a] and ODE sampling [Song et al., 2020b], randomness is only present at the start of the sampling process (akin to VAEs and GANs). In contrast, in DDPM [Ho et al., 2020] and SDE sampling [Song et al., 2020b], randomness is introduced at every step throughout the sampling process. Generative uncertainty is then defined as the variability of the posterior predictive:
@@ 3.2 Last-Layer Laplace Approximation
State-of-the-art diffusion models are extremely large (100M to 1B+ parameters) and can take weeks to train. Consequently, the computational overhead of performing Bayesian inference on such large models is of significant concern. To address this, we adopt the Laplace approximation [MacKay, 1992a, Shun and McCullagh, 1995] to approximate the posterior $q(\theta|\mathcal{D})$. The Laplace approximation is among the most computationally efficient approximate inference methods while still offering competitive performance [Daxberger et al., 2021a]. Moreover, a particularly appealing feature of the Laplace approximation is that it can be applied post-hoc to any diffusion model. We leverage this property in Section 4, where we apply it to a variety of popular diffusion and flow-matching models.
@@
We note that the use of last-layer Laplace approximation for diffusion models has been previously proposed in BayesDiff [Kou et al., 2024]. While our implementation of the Laplace approximation closely follows theirs, there are significant differences in how we utilize the approximate posterior, $q(\theta|\mathcal{D})$. Specifically, in our approach, we use it within the traditional Bayesian framework (Eq. 3) to sample new diffusion model parameters, leaving the diffusion sampling process, $g_{\theta}$, unchanged. In contrast, BayesDiff resamples new weights from $q(\theta|\mathcal{D})$ at every diffusion sampling step $t$, which necessitates substantial modifications to the diffusion sampling process through their variance propagation approach. We later demonstrate in Section 4 that modifications such as variance propagation are unnecessary for obtaining Bayesian generative uncertainty and staying closer to the traditional Bayesian setting leads to the best empirical performance.
3.3 Semantic Likelihood
where $\sigma^{2}$ represents the observation noise.
However, as we will demonstrate in Section 4, this likelihood leads to non-informative estimates of generative uncertainty (Eq. 4). The primary issue is that the sample space of natural images is high-dimensional (i.e., $|\mathcal{X}|=HWC$). Consequently, placing the likelihood directly in the sample space causes the variability of the posterior predictive distribution to be based on pixel-level differences. This is problematic because it is well-known that two images can appear nearly identical to the human eye while exhibiting a large $L_{2}$-norm difference in pixel space $\mathcal{X}$ (see, for example, the literature on adversarial examples [Szegedy, 2013]). To get around this, we propose to map the generated samples to a ‘semantic’ latent space, $\mathcal{S}$, via a pre-trained feature extractor, $c_{\phi}:\mathcal{X}\rightarrow\mathcal{S}$ (e.g., an inception-net [Szegedy et al., 2016] or a CLIP encoder [Radford et al., 2021]). The resulting semantic likelihood has the form
where $\mathbf{e}(\mathbf{x})\in\mathcal{S}$ is the (random) vector of semantic features.
By combining the (last-layer) Laplace approximate posterior and the semantic likelihood, we can now approximate the posterior predictive (Eq. 3) as
where $M$ denotes the number of Monte Carlo samples. Additionally, we approximate the posterior predictive with a single Gaussian via moment matching here, a common practice in Bayesian neural networks for regression problems [Lakshminarayanan et al., 2017, Antorán et al., 2020].
@@ 9       $\boldsymbol{e}_{m}=c_{\phi}(\hat{\boldsymbol{x}}_{m})$
13Compute $p(\boldsymbol{x}|\boldsymbol{z},\mathcal{D})$ using $\{\boldsymbol{e}_{m}\}_{m=0}^{M}$ (Eq. 7)
@@
3.4 Epistemic Uncertainty
Uncertainty is commonly decomposed into two components: aleatoric and epistemic [Hüllermeier and Waegeman, 2021, Smith et al., 2024]. Aleatoric uncertainty represents the irreducible uncertainty inherent in the data-generating process, while epistemic uncertainty arises from observing only a limited amount of training data. In our framework, we fix the observation noise in the semantic likelihood (Eq. 7) to a small constant value (e.g., $\sigma=0.001$). As a result, the generative uncertainty we capture is primarily epistemic, reflecting uncertainty about the model parameters $\theta$ due to limited training data via $q(\theta|\mathcal{D})$. Since the parameters $\phi$ of the semantic feature extractor $c_{\phi}$ are kept fixed in the semantic likelihood, the resulting generative uncertainty $u(\boldsymbol{z})$ continues to reflect the epistemic uncertainty of the diffusion model parameters $\theta$. Extending our framework to capture the aleatoric uncertainty of a generative process presents an interesting avenue for future research.
@@ 4 Experiments
In our experiments, we demonstrate that generative uncertainty is an effective method for detecting poor samples in diffusion models (Section 4.1). We also discuss the sampling overhead introduced by our Bayesian approach and show that it can be effectively minimized (Section 4.2). Finally, we extend our Bayesian framework beyond diffusion by applying it to detect low-quality samples in a (latent) flow matching model (Appendix A.7). Our code is available at https://github.com/metodj/DIFF-UQ.

### v2

@@ 1 Introduction
Diffusion (and flow-matching) models [Sohl-Dickstein et al., 2015, Song et al., 2021a, c, Lipman et al., 2023] have recently pushed the boundaries of generative modeling due to their strong theoretical underpinnings and scalability. Across various domains, they have enabled the generation of increasingly realistic samples [Rombach et al., 2022, Esser et al., 2024, Li et al., 2024]. Despite the impressive progress, state-of-the-art models can still generate low quality images that contain artefacts and fail to align with the provided conditioning information. This poses a challenge for deploying diffusion models, as it can lead to a poor user experience by requiring multiple generations to manually find an artefact-free sample.
@@
    We show that our generative uncertainty strongly outperforms previous uncertainty-based approaches for filtering out poor samples [Kou et al., 2024, De Vita and Belagiannis, 2025] (Section 4.2). Additionally, we achieve competitive performance with non-uncertainty-based methods, such as realism score [Kynkäänniemi et al., 2019] and rarity score [Han et al., 2023], while also highlighting the complementary benefits of uncertainty (Appendix C.5).
@@
    We propose effective strategies to reduce the sampling overhead of Bayesian uncertainty (Section 4.3) and demonstrate the applicability of our framework beyond diffusion models by applying it to a (latent) flow matching model (Section C.7).
@@ 2.1 Generative Modeling
Modern deep generative models like variational autoencoders (VAEs) [Kingma et al., 2014], generative adversarial networks (GANs) [Goodfellow et al., 2014], and diffusion models differ in their exact probabilistic frameworks and training schemes, yet share a common sampling recipe: start with random noise and transform it into a new data sample [Tomczak, 2022]. Specifically, let $\boldsymbol{x}\in\mathcal{X}$ denote a data sample and $\boldsymbol{z}\in\mathcal{Z}$ an initial noise. A new sample is generated by:
@@
where $p(\mathbf{z})$ is an initial noise (prior) distribution, typically a standard Gaussian $\mathcal{N}(0,I)$, and $g_{\theta}:\mathcal{Z}\rightarrow\mathcal{X}$ is a generator function with model parameters $\theta\in\mathbb{R}^{P}$. Here and throughout the paper, we use $\boldsymbol{z}$ (with a slight abuse of notation) to denote the entire randomness involved in the sampling process. [^1] This distinction matters for diffusion models. In DDIM [Song et al., 2021a] and ODE sampling [Song et al., 2021c], randomness is only present at the start of the sampling process (akin to VAEs and GANs). In contrast, in DDPM [Ho et al., 2020] and SDE sampling [Song et al., 2021c], randomness is introduced at every step throughout the sampling process.
@@
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ and $\{\beta_{s}\}_{s=1}^{T}$, with $\beta_{S}\in(0,1)$, is a noise schedule chosen such that $\boldsymbol{x}_{T}\sim\mathcal{N}(0,I)$ (approximately). In the backward process, a denoising network $f_{\theta}$ is learned via a simplified regression objective (among various possible parameterizations, see Song et al. [2021c] or Karras et al. [2022]):
$\!\!\!\mathcal{L}(\theta)=\mathbb{E}_{t,\boldsymbol{x}_{0},\boldsymbol{\epsilon}}\!\left[\big{|}\big{|}f_{\theta}(\sqrt{\bar{\alpha}_{t}}\boldsymbol{x}_{0}\!+\!\sqrt{1\!-\!\bar{\alpha_{t}}}\boldsymbol{\epsilon},t)-\boldsymbol{\epsilon}\big{|}\big{|}_{2}^{2}\right]\>.$ (1)
After training, diffusion models generate new samples via a generator function, $g_{\hat{\theta}}$, which consists of sequentially applying the learned denoiser, $f_{\hat{\theta}}$, and following specific transition rules from samplers such as DDPM [Ho et al., 2020] or DDIM [Song et al., 2021a].
Figure 1: Generative uncertainty. Illustration of how we compute generative uncertainty for a fixed random noise $\boldsymbol{z}$. For a diffusion model $g_{\theta}$, we draw $M$ parameter sets $\{\theta_{1},…,\theta_{M}\}$ from the approximate posterior $q(\theta|D)$ over diffusion model’s parameters (here: last-layer Laplace with $M=3$). Each model $g_{\theta_{m}}$ maps $\boldsymbol{z}$ to an image $\hat{\boldsymbol{x}}_{m}$. Embedding all $\hat{\boldsymbol{x}}_{m}$ with a frozen encoder $c_{\phi}$ (e.g., CLIP) gives $M$ semantic feature vectors $\boldsymbol{e}_{m}$; the variability (e.g., entropy) of these vectors is the final uncertainty $u(\boldsymbol{z})$. Low-uncertainty (left) corresponds to consistent, high-quality images, whereas high-uncertainty (right) reveals model disagreement and poor, discordant outputs.
@@ 2.2 Bayesian Deep Learning
Bayesian neural networks (BNNs) go beyond point predictions and allow for principled uncertainty quantification [Neal, 1995, Kendall and Gal, 2017, Jospin et al., 2022]. Let $h_{\psi}:\mathcal{X}\rightarrow\mathcal{Y}$ denote a predictive model with parameters $\psi\in\mathbb{R}^{O}$ and $\mathcal{D}=\{(\boldsymbol{x}_{n},\boldsymbol{y}_{n})\}_{n=1}^{N}$ denote training data. Instead of finding a single fixed set of parameters, $\hat{\psi}=\arg\max\mathcal{L}(\psi;\mathcal{D})$, that maximizes a chosen objective function $\mathcal{L}$, BNNs specify a prior $p(\psi)$ over model parameters and define a likelihood $p(\mathbf{y}|h_{\psi}(\boldsymbol{x}))$, which together yield a posterior distribution via Bayes rule: $p(\psi|\mathcal{D})\propto p(\psi)\prod_{n=1}^{N}p(\boldsymbol{y}_{n}|h_{\psi}(\boldsymbol{x}_{n}))$. Under this Bayesian view, a predictive model for a new test point $\boldsymbol{x}_{*}$ is then obtained via the posterior predictive distribution:
@@
For large models, finding the exact posterior distribution is computationally intractable, hence an approximate posterior $q(\psi|\mathcal{D})$ is used instead. Popular approaches for approximate inference include deep ensembles [Lakshminarayanan et al., 2017, Wilson and Izmailov, 2020], variational inference [Blundell et al., 2015, Zhang et al., 2018], SWAG [Mandt et al., 2017, Maddox et al., 2019], and Laplace approximation [Daxberger et al., 2021a]. Moreover, to alleviate computational overhead, it is common to give a ‘Bayesian treatment’ only to a subset of parameters [Kristiadi et al., 2020, Daxberger et al., 2021b, Sharma et al., 2023]. Finally, the intractable expectation integral in the posterior predictive is approximated via Monte-Carlo (MC) sampling:
@@ 3.1 Generative Uncertainty
Here, as before in Section 2.1 , we use $\boldsymbol{z}$ to denote the entire randomness involved in the diffusion sampling process. Generative uncertainty is then defined as the variability of the posterior predictive:
@@ 3.2 Last-Layer Laplace Approximation
State-of-the-art diffusion models are extremely large (100M to 1B+ parameters) and can take weeks to train. Consequently, the computational overhead of performing Bayesian inference on such large models is of significant concern. For instance, while deep ensembles are a convenient and popular approach for predictive models [Lakshminarayanan et al., 2017], the sheer size of diffusion models renders naive ensembling infeasible. To address this, we adopt the Laplace approximation [MacKay, 1992a, Shun and McCullagh, 1995] to find the approximate posterior $q(\theta|\mathcal{D})$. The Laplace approximation is among the most computationally efficient approximate inference methods while still offering competitive performance [Daxberger et al., 2021a]. Moreover, a particularly appealing feature of the Laplace approximation is that it can be applied post-hoc to any diffusion model. We leverage this property in Section 4, where we apply it to a variety of popular diffusion and flow-matching models.
@@
Note that for the Laplace approximation to be theoretically valid, the loss function should correspond to (log-)likelihood and (log-)prior terms. While the diffusion loss can be interpreted as a log-likelihood up to an additive constant—due to its role as a surrogate for the KL divergence between the noising and denoising processes—this interpretation holds only under appropriate weighting of the loss terms across different time steps [Song et al., 2021b]. Consequently, applying the Laplace approximation directly, without such reweighting, is not fully theoretically justified. Despite this, we find in our experiments (Section 4.2) that applying the Laplace approximation directly to the diffusion losses used in practice (without modifying the loss weighting) still yields meaningful uncertainty estimates that align well with the visual quality of generated samples. That said, we believe that better understanding the theoretical requirements for applying approximate Bayesian inference techniques like Laplace in modern generative models like diffusion remains an important direction for future work that could lead to even more informative uncertainty estimates.
It is also worth noting that the use of last-layer Laplace approximation for diffusion models has been previously proposed in BayesDiff [Kou et al., 2024]. While our implementation of the Laplace approximation closely follows theirs, there are significant differences in how we utilize the approximate posterior, $q(\theta|\mathcal{D})$. Specifically, in our approach, we use it within the traditional Bayesian framework (Eq. 3) to sample new diffusion model parameters, leaving the diffusion sampling process, $g_{\theta}$, unchanged. In contrast, BayesDiff resamples new weights from $q(\theta|\mathcal{D})$ at every diffusion sampling step $t$, which necessitates substantial modifications to the diffusion sampling process through their variance propagation approach. We later demonstrate in Section 4.2 that modifications such as variance propagation are unnecessary for obtaining Bayesian generative uncertainty and staying closer to the traditional Bayesian setting leads to the best empirical performance.
@@ 9       $\boldsymbol{e}_{m}=c_{\phi}(\hat{\boldsymbol{x}}_{m})$
13Compute $p(\boldsymbol{x}|\boldsymbol{z},\mathcal{D})$ using $\{\boldsymbol{e}_{m}\}_{m=0}^{M}$ (Eq. 3.3)
@@
3.3 Semantic Likelihood
where $\sigma^{2}$ represents the observation noise.
Figure 2: Illustration of how generative uncertainty can be effective for filtering our poor generations on a simple 2D Gaussian dataset. Left: training data consisting of 25 separate Gaussian modes. Middle: generated samples using a trained diffusion model and a DDPM sampler ($T=1000$). Right: same set of generated samples after removing 50% of generations with the highest estimated generative uncertainty.
However, as we will demonstrate in Section 4, this likelihood leads to non-informative estimates of generative uncertainty (Eq. 4). The primary issue is that the sample space of natural images is high-dimensional (i.e., $|\mathcal{X}|=HWC$). Consequently, placing the likelihood directly in the sample space causes the variability of the posterior predictive distribution to be based on pixel-level differences. This is problematic because it is well-known that two images can appear nearly identical to the human eye while exhibiting a large $L_{2}$-norm difference in pixel space $\mathcal{X}$ (see, for example, the literature on adversarial examples [Szegedy et al., 2013]). To get around this, we propose to map the generated samples to a ‘semantic’ latent space, $\mathcal{S}$, via a pre-trained feature extractor, $c_{\phi}:\mathcal{X}\rightarrow\mathcal{S}$ (e.g., an inception-net [Szegedy et al., 2016] or a CLIP encoder [Radford et al., 2021]). The resulting semantic likelihood has the form
where $\mathbf{e}(\mathbf{x})\in\mathcal{S}$ is the (random) vector of semantic features. By combining the (last-layer) Laplace approximate posterior and the semantic likelihood, we can now approximate the posterior predictive (Eq. 3) as
where $M$ denotes the number of Monte Carlo samples. Additionally, we approximate the posterior predictive with a single Gaussian via moment matching here, a common practice in Bayesian neural networks for regression problems [Lakshminarayanan et al., 2017, Antorán et al., 2020]. A more detailed derivation is provided in Appendix B.1.
@@ 4 Experiments
In our experiments, we begin by demonstrating that generative uncertainty serves as an effective method for identifying poor samples in diffusion models, using a simple synthetic dataset (Section 4.1). We then show that our proposed approximations (Sections 3.2 and 3.3) enable the estimation of generative uncertainty in large, modern diffusion models applied to high-dimensional natural images (Section 4.2). Additionally, we discuss the sampling overhead introduced by our Bayesian approach and show that it can be effectively reduced (Section 4.3). Finally, we extend our Bayesian framework beyond diffusion by applying it to detect low-quality samples in a (latent) flow matching model (Appendix C.7). Our code is available at https://github.com/metodj/DIFF-UQ.
4.1 Toy Demonstration
To illustrate the potential of generative uncertainty for detecting poor generations, we adopt the setting from Aithal et al. [2024]. Specifically, we use a 2D synthetic dataset with 25 distinct modes (Figure 2, left) to train a small diffusion model. We then generate 50K samples using a DDPM sampler (Figure 2, middle). While the generated samples cover the 25 modes well, many ‘hallucinated’ samples also appear between the modes of the training data. Following Aithal et al. [2024], we consider such samples to be poor generations, as they are highly unlikely under the true data-generating distribution.
Next, we train an ensemble of diffusion models ($M=5$) and use it to estimate the generative uncertainty of each of the 50K generated samples. We then filter out the 50% of samples with the highest estimated uncertainty and plot the remaining ones in Figure 2, right. As shown in the plot, this uncertainty-based filtering effectively removes all poor generations between modes, indicating that generative uncertainty can serve as a reliable indicator of sample quality. Note that in this simple toy setting, we use neither the Laplace approximation (relying instead on a diffusion deep ensemble) nor the semantic likelihood. In the following section, we show how both can be employed to extend generative uncertainty estimation to the more realistic setting of natural images.